# CaTS-Bench: Can Language Models Describe Numeric Time Series?

### v1

@@ 3.1 Data Curation
# Samples Generated 20k 4.4k 3.2k 764 598 756 5.5k 732 2.1k 544 551 835
# Train Samples 16k 3.5k 2.6k 611 478 604 4.4k 585 1.7k 435 440 668
Avg. Sample Length 29.1 65.3 21.2 76.8 11.6 5.9 75.8 9.5 12.2 12.2 22.4 7.3
# Test Samples 4k 886 646 153 120 152 1.1k 147 422 109 111 167
Avg. Sample Length 26.1 66.0 21.2 76.9 5.0 3.6 73.0 8.7 5.5 11.8 8.1 7.5
# Human-revisited Samples 579 0 0 153 120 0 0 0 0 109 0 167
Avg. Sample Length 25.7 - - 76.9 5.0 - - - - 11.8 - 7.5
To prevent information leakage, we partition each source dataset temporally before generating the samples. Specifically, the first $80\%$ is used for generating training samples, whereas the last $20\%$ is reserved exclusively for generating test samples. Random window cropping is applied separately to the training and test partitions. This strategy ensures that the model is evaluated on future, unseen data relative to the training set. The actual benchmark samples consist of the test split resulting from this process. We leave the training split of the data for optional training. Our final semi-synthetic dataset version contains $20k$ examples, split into roughly $16k$ training samples and $4k$ test samples. Detailed statistics and source of our data are reported in Table 2.
Human-Revisited Subset.
We also release a curated subset of test captions that have been revisited by humans. These captions were first sampled from multiple LLM candidates (Gemini 2.0 Flash, GPT-4o, Gemma 27B, and Llama 90B) using the above data pipeline, and then carefully refined by the authors to eliminate factual errors, speculative statements, and redundant phrasing. Drawn from the domains of agriculture, crime, demography, and Walmart sales, this subset provides high-fidelity, human-styled references that complement the larger benchmark.
3.2 Quality Validation of Semi-synthetic Captions
To ensure the quality of CaTS-Bench, we conducted a series of comprehensive verification studies addressing the core concern of semi-synthetic data: whether captions generated by the oracle model (Gemini 2.0 Flash) are factual, unbiased, and linguistically diverse. These analyses demonstrate that semi-synthetic captions in CaTS-Bench provide high-quality references and stable benchmarks for TSC, and thus are a sufficient proxy for human-written descriptions in practical scenarios. We verified caption quality through three complementary studies below (with full details in Appendix H).
Manual Validation. We manually checked $\sim$$2.9k$ captions (72.572.5% of the semi-synthetic test benchmark) across statistical claims (min, max, mean, STD) and trend descriptors (up/downward, stable, fluctuating). Accuracy exceeded 98.6% on average across all categories (Table 9) which confirms that captions faithfully reflect underlying series properties.
Human Detectability Study. In a blind test with $35$ participants, subjects attempted to distinguish our captions from those written by humans. Accuracy was near random at $41.1\%$, suggesting that our captions are indistinguishable from human-authored ones and no evidence of oracle-specific bias.
Diversity and Bias Analysis. Captions consistently drew from a wide variety of statistical and temporal descriptors (Table 12), and embedding-based similarity analysis across nine embedding models revealed minimal template reliance. Pairs of captions that were almost semantically identical, measured as embedding cosine similarity >0.95>0.95, were rare, averaging $2.3\%$ of occurrences (Table 13). Comparisons with human captions ( H.4.4) indicate that Gemini’s outputs are stylistically intermixed with human text, while N-gram analysis (H.4.2) confirms high lexical diversity.
3.3 Time Series Captioning
TSC requires generating a detailed, coherent narrative that highlights the key characteristics of a given time series. During evaluation, each model is presented with a standardized multi-part prompt that combines four elements: the numeric series itself, embedded as raw time-indexed values (e.g., [25.3, 26.1, 26.8, …]); contextual metadata such as measurement units, data source, sampling interval, and domain tags (e.g., “Hourly temperature readings from Rome, May 2000”), which excludes explicit statistics like mean or maximum since the model must infer them; a visual input in the form of a line-plot image that allows vision-language models to ground their descriptions in visual trend cues; and a fixed-format instruction template containing the directive for caption generation (see Appendix N.2). By standardizing this multi-part prompt, we evaluate models on their ability to recognize numeric trends (e.g., rising or falling segments, peaks, and troughs), integrate metadata cues, and utilize visual features to produce context-aware captions.
3.4 Q&A Multiple-Choice Tasks
We introduce a suite of multiple-choice Q&A tasks designed to probe different reasoning skills in time series understanding. All tasks are automatically derived from the same source data used for captioning, with questions generated from task-specific, fixed templates (see Appendix J.1 for examples). To increase difficulty, an initial pool of $4k$ questions per type was filtered by removing those correctly answered by Qwen 2.5 Omni. Appendix J.2 shows that this filtering produces genuinely harder questions, rather than reflecting Qwen-specific weaknesses only. Ambiguous Time Series Matching questions were manually checked to ensure a single correct answer. From the remaining $7k$ challenging questions, a random subset of $460$ was sampled as the final test set, including $100$ each for time series matching, caption matching, and plot matching, and $40$ each for amplitude, peak, mean, and variance comparison tasks. Question types are described below.
Time Series Matching. Given a caption, the model must retrieve the correct time series from distractor candidates created via shuffling, temporal reversal, and Gaussian noise. These perturbations prevent simple numeric lookup and require alignment with both values and trends (see J.3 for details).
Caption Matching. Given a time series, the model must select the correct caption from distractors composed of random captions and perturbed variants of the ground truth (see Appendix N.5N.6). This isolates caption understanding from free-form generation.
Plot Matching. Given a caption and its numeric series, the model must select the correct line plot from the candidates, testing visual grounding and the ability to link language with visual patterns.
Time Series Comparison.
Given two time series, select the correct comparative statement from a pair of options (e.g., “Series A peaks earlier than Series B” or “Series B has a higher volatility than Series A”). This task challenges models to perform temporal and statistical comparison, a setting where many language models currently struggle (Merrill et al., 2024).
@@ 3.5 Evaluation Metrics
To comprehensively evaluate model-generated captions against the ground truth in TSC, we employ a diverse set of metrics that target linguistic quality, statistical inference, and numeric fidelity. For Q&A, we adopt accuracy as the evaluation metric, as each question is designed to have a single correct answer. Below, we describe each metric used for TSC in our evaluation framework.
Standard Linguistic Metrics. We assess caption similarity using standard NLP metrics, including DeBERTa Score (Zhang* et al., 2020), BLEU (Papineni et al., 2002), ROUGE-L (Chin-Yew, 2004), METEOR (Banerjee & Lavie, 2005), and SimCSE (Gao et al., 2021; Liu et al., 2019). Together, these metrics capture both surface-level linguistic overlap and deeper semantic similarity. This ensures that evaluation does not merely reflect stylistic resemblance but instead rewards accurate semantics of the underlying time series phenomena. Refer to Appendix F for more details.
numeric Fidelity Metrics. Since TSC involves reporting exact or approximate numeric values, we introduce two tailored metrics to quantify numeric accuracy, both bounded within $[0,1]$. The choice of the $5\%$ tolerance is discussed in Appendix F.2.
@@
    Statistical Inference Accuracy. While models are explicitly prompted to discuss descriptive statistics, they demonstrate varying abilities to accurately infer and verbalize statistics such as the mean, standard deviation, minimum, and maximum based on the raw time series and metadata. To evaluate this behavior, we report the percentage of captions in which these statistics are mentioned and fall within a $5\%$ relative error, using offline-computed true values. Importantly, captions are not penalized for omitting statistics; only wrongly reported values are considered errors. This metric primarily measures hallucination, favoring omission over incorrect numeric claims.
@@
    Numeric Score. For each ground truth caption, we extract all numeric values (excluding time-related ones like year or month) and search for the closest numeric value in the generated caption. A match is recorded if the closest value is within a $5\%$ relative tolerance. We compute Accuracy (mean of $1-\min\{\text{relative\_error},\text{tolerance}\})$ over all matched numbers), Recall (fraction of ground truth numbers matched), and a Final Score as a weighted combination: $\lambda_{A}\cdot\text{Accuracy}+\lambda_{R}\cdot\text{Recall}$, with $\lambda_{A}=0.3$ and $\lambda_{R}=0.7$ to emphasize recall over precision, as omitting critical numbers is more severe than minor numeric rounding imprecisions. While the previous metric targets numeric hallucinations, this one focuses on penalizing captions that omit numeric details.
@@ 4 Experiments
We evaluate a broad range of VLMs on CaTS-Bench, covering both proprietary and open-source models, with the latter also tested after finetuning on our captioning training set (details in Appendix D). For TSC, we additionally consider a program-aided (PAL) model (Gao et al., 2023). All models are prompted with the same template-based format to ensure fair comparison, avoiding task- or architecture-specific prompt engineering. Appendix E provides the full model list and a description of PAL, while Appendix O outlines the human baseline that participated in our Q&A evaluation.
4.1 Time Series Captioning
@@
To ensure fair comparison across the domains, we report macro-averaged scores for each metric, mitigating sample size imbalances, as some domains contain more data, and preventing any domain from disproportionately influencing the results. We benchmark leading VLMs on TSC using the semi-synthetic and human-revisited captions separately as ground truth. Selected results are shown in Tables 3 and 4, with complete results in the Appendix G.
Semi-synthetic (SS) Captions as Ground Truth.
Table 3: Selected evaluation results of generated captions against human-revisited (HR) and semi-synthetic ($S$) ground truths. Bolded and underlined scores denote first and second places.
Category Model DeBERTa F1 SimCSE BLEU ROUGE-L METEOR Numeric
HR $S$ HR $S$ HR $S$ HR $S$ HR $S$ HR $S$
Proprietary
Gemini 2.0 Flash 0.665 0.688 0.856 0.858 0.079 0.137 0.248 0.318 0.221 0.279 0.634 0.677
Gemini 2.5 Pro 0.657 0.668 0.857 0.845 0.069 0.088 0.236 0.267 0.247 0.284 0.681 0.714
Claude 3 Haiku 0.658 0.682 0.853 0.856 0.064 0.112 0.241 0.291 0.236 0.300 0.601 0.623
GPT-4o 0.661 0.681 0.863 0.865 0.071 0.112 0.233 0.284 0.236 0.296 0.627 0.644
Pretrained
InternVL 2.5 38b 0.664 0.688 0.871 0.868 0.072 0.129 0.244 0.305 0.255 0.331 0.659 0.685
LLaVA v1.6 0.627 0.650 0.824 0.820 0.052 0.086 0.215 0.259 0.233 0.287 0.455 0.517
LLaVA v1.6 34b 0.639 0.655 0.821 0.825 0.060 0.094 0.221 0.265 0.232 0.285 0.547 0.560
Idefics 2 0.602 0.604 0.784 0.698 0.024 0.040 0.192 0.226 0.140 0.162 0.424 0.455
SmolVLM 0.592 0.594 0.755 0.693 0.027 0.044 0.194 0.224 0.154 0.178 0.431 0.474
QwenVL 0.619 0.643 0.821 0.890 0.049 0.082 0.209 0.249 0.214 0.261 0.445 0.504
QwenVL PAL 0.664 0.685 0.864 0.843 0.066 0.108 0.237 0.292 0.226 0.282 0.564 0.613
Llama 3.2 V 0.653 0.671 0.852 0.850 0.072 0.118 0.239 0.290 0.252 0.315 0.650 0.685
Gemma 3 27b 0.648 0.667 0.863 0.863 0.065 0.085 0.222 0.263 0.257 0.309 0.641 0.668
Finetuned
LLaVA v1.6 0.712 0.758 0.896 0.907 0.134 0.285 0.312 0.445 0.300 0.441 0.693 0.732
Idefics 2 0.711 0.759 0.894 0.908 0.132 0.290 0.309 0.452 0.298 0.437 0.691 0.733
InternVL-2.5 8b 0.638 0.655 0.817 0.809 0.053 0.088 0.215 0.259 0.229 0.282 0.582 0.594
QwenVL 0.703 0.643 0.892 0.790 0.126 0.082 0.302 0.249 0.297 0.260 0.683 0.504
SmolVLM 0.604 0.613 0.817 0.781 0.051 0.091 0.228 0.269 0.220 0.265 0.556 0.643
Our experiments show that finetuning substantially improves performance across most metrics. Proprietary models such as GPT-4o and Gemini generally outperform Claude. Among open-source models, finetuned Idefics 2 and LLaVA v1.6 Mistral achieve strong gains, in some cases surpassing proprietary baselines, underscoring the effectiveness of finetuning for both linguistic quality and numeric precision. QwenVL PAL shows marked improvements over standard QwenVL and even takes the lead on statistical inference metrics (as shown in Table 4), highlighting code execution as a practical enhancement for tasks where numbers matter.
Table 4: Representative statistical inference scores under ground truths. E.g., Mean indicates statistical inference of the series mean. Bolded and underlined scores denote first and second places.
Category Model Mean Max Min
HR $S$ HR $S$ HR $S$
Proprietary
Gemini 2.0 Flash 0.536 0.651 0.982 0.985 0.936 0.917
Gemini 2.5 Pro Prev. 0.323 0.494 0.987 0.994 0.977 0.971
Claude 3 Haiku 0.833 0.693 0.980 0.977 0.934 0.898
GPT-4o 0.817 0.700 0.992 0.990 0.938 0.921
Pretrained
InternVL 2.5 38b 0.858 0.784 0.982 0.966 0.930 0.887
LLaVA v1.6 Mistral 0.667 0.644 0.871 0.864 0.751 0.743
LLaVA v1.6 34b 0.410 0.445 0.817 0.843 0.727 0.698
Idefics 2 0.806 0.616 0.891 0.903 0.840 0.806
QwenVL 0.656 0.565 0.795 0.822 0.678 0.657
QwenVL PAL 0.973 0.903 0.985 0.980 0.978 0.942
Llama 3.2 Vision 0.467 0.594 0.956 0.952 0.895 0.877
Gemma 3 27b 0.734 0.694 0.978 0.968 0.904 0.864
Finetuned
LLaVA v1.6 Mistral 0.928 0.828 0.987 0.976 0.981 0.926
Idefics 2 0.958 0.885 0.988 0.985 0.967 0.927
InternVL 2.5 (8b) 0.750 0.597 0.830 0.904 0.734 0.779
QwenVL 0.952 0.565 0.973 0.822 0.963 0.657
SmolVLM 0.640 0.590 0.914 0.898 0.772 0.777
Given the semi-synthetic nature of ground truths in this experiment, we assessed the robustness of evaluation along two axes. First, to account for the stochasticity of LLM outputs, we repeated inference three times on $\sim 600$ test samples across five representative models; variance was vanishingly small (often $10^{-6}$; Appendix H.5), confirming that our single-run results are stable and reliable. Second, to test sensitivity to linguistic style, we paraphrased a subset of ground truth captions using multiple architecturally distinct LLMs while strictly preserving all factual content and numeric details, generating variants of ground truths differing only by linguistic style. The paraphrasing prompt is provided in Appendix N.3. Re-evaluating baseline outputs against these paraphrased captions as ground truth yielded model performance rankings largely consistent with those based on the original Gemini captions, with a mean Spearman Correlation of $0.9266$ and metric-specific correlations shown in Table 11 (full discussion in  H.3). These results corroborate that our evaluation framework is stable and reliably gauges caption quality rather than biased surface-level stylistic alignment.
Human-revisited (HR) Captions as Ground Truth.
We repeat the evaluation using human-revisited captions as ground truth, further confirming the benefits of finetuning. Open-source models like Idefics 2 and LLaVA v1.6 Mistral gain substantially in text quality and numeric accuracy, often surpassing proprietary baselines on linguistic metrics and nearing them on numeric ones. Proprietary models such as GPT-4o and Gemini still lead on some language-focused metrics, but their advantage shrinks when finetuned open-source models are included. Meanwhile, the PAL model excels in statistical inference thanks to code execution. Overall, these results confirm that finetuning not only enhances average performance but also improves numeric reliability, positioning open-source models as strong contenders when paired with targeted adaptation.
4.2 Q&A Tasks
Figure 3 summarizes model performance on our Q&A tasks, while Table 17 provides detailed results. Performance is highly variable, and even proprietary models occasionally fail to exceed random chance on some tasks.

### v2

@@ 3.1 Base Data Curation
# Triplets Generated 20k 4.4k 3.2k 764 598 756 5.5k 732 2.1k 544 551 835
# Human Test Samples 1746 0 323 153 120 152 300 100 211 109 111 167
Average TS Length 26.8 - 21.5 76.9 5.0 3.6 71.8 8.8 5.5 11.8 8.1 7.5
# Synth. Train Samples 16k 3.5k 2.6k 611 478 604 4.4k 585 1.7k 435 440 668
Average TS Length 29.1 65.3 21.2 76.8 11.6 5.9 75.8 9.5 12.2 12.2 22.4 7.3
# Synth. Test Samples 4k 886 646 153 120 152 1.1k 147 422 109 111 167
Average TS Length 26.1 66.0 21.2 76.9 5.0 3.6 73.0 8.7 5.5 11.8 8.1 7.5
3.2 Data Generation Streams
Using the base data triplets as a foundation, we employ two distinct streams to generate natural language descriptions: one focused on HR evaluation and the other on scalable training.
Human-Rewritten Gold Standard.
To mitigate the potential oracle bias and stylistic homogeneity inherent in single-model benchmarks, we curated a primary evaluation set of $1746$ human-rewritten captions. To ensure broad linguistic coverage and minimize model-specific artifacts, we utilized a multi-stage curation process: we initially generated candidate drafts from four distinct LLMs (Gemini 2.0 Flash, GPT-4o, Gemma 27B, and Llama 3 70B). Using these diverse drafts as structural scaffolds, we performed extensive rewriting and factual verification (see Appendix K for details and results of the editing process and Appendix G.4 for diversity analysis). Across all domains, human rewriting resulted in substantial linguistic revision (average WER = $0.60$) while maintaining strong numeric overlap (average Numeric Jaccard = $0.74$). This approach allowed us to move beyond simple post-editing, resulting in captions that are factually grounded, linguistically diverse, and enriched with the interpretive nuances that purely automated systems typically miss. These HR descriptions serve as the gold standard for CaTS-Bench, providing a rigorous test of a model’s ability to generate high-fidelity, context-aware narratives. Beyond evaluation, these captions reflect how humans summarize time series for rapid comprehension.
Synthetic Augmentation Pipeline.
To facilitate large-scale training, we utilize a scalable pipeline to generate additional reference captions. We query an oracle LLM (Gemini 2.0 Flash) with a structured prompt comprising the full numeric sequence, enriched metadata with precomputed statistics, and the corresponding line plot. The oracle is strictly instructed to remain within the bounds of the provided data, ensuring that the generated captions are factually grounded. (see Appendix L.1 for prompt example). We emphasize that the use of a single oracle model is intended as a proof of concept; different or multiple oracles can be used. While the synthetic data generation primarily serves as a large-scale training resource, we also utilize it to establish a dual-evaluation framework. We generate a synthetic corpus of approximately 2020k captions. Of these, 1616k are used for training, while the full synthetic test split contains roughly 44k captions. For evaluation, however, we restrict attention to a subset of $1746$ synthetic captions that exactly correspond to the HR evaluation set in underlying time series. This dual-track allows us to compare synthetic and HR captions on identical samples, revealing where model-generated descriptions diverge from human-written ones.
Figure 2: Overview of CaTS-Bench’s dual-stream data generation pipeline. A time series window is cropped, metadata is attached, and a human or LLM writes a reference caption. See Appendix N for examples and Appendix G for the quality verification protocol.
3.3 Synthetic Data Quality Validation
A key concern when using synthetic captions for training is whether they are factually accurate, linguistically diverse, and free from artifacts that hinder generalization to human-authored text. To validate our synthetic data pipeline, we conduct three complementary analyses, with full protocols and results reported in Appendix G.
First, we perform a manual factual validation over approximately $2.9k$ synthetic captions, verifying statistical claims and trend descriptions against the underlying numeric series. This analysis shows that over $98.6\%$ of captions are factually correct, indicating minimal numeric hallucination. Second, we conduct a human indistinguishability study with $35$ participants, who are asked to distinguish synthetic captions from human-written ones in a blind setting. The resulting detection accuracy is $41.1\%$, nearing random chance, suggesting that the synthetic captions are stylistically comparable to human-written descriptions. Finally, we assess linguistic diversity to rule out template-driven generation. Embedding-based similarity analysis shows that near-duplicate captions (cosine similarity >0.95>0.95) occur in only $2.3\%$ of cases, and complementary N-gram and vocabulary analyses confirm broad lexical coverage comparable to human text. These results show that the synthetic captions provide high-fidelity training signals that closely approximate human descriptions.
3.4 Task 1: Time Series Captioning
The primary objective of TSC is the generation of a linguistically coherent narrative describing the salient temporal characteristics of a numeric series. During evaluation, models are presented with a multimodal prompt comprising four components: i) numeric series consisting of raw, time-indexed values that require direct numeric reasoning; ii) contextual metadata providing semantic background (e.g., measurement units and location), while explicitly excluding any statistical summaries of that series to ensure models perform genuine statistical inference; iii) a visual representation via a line-plot image that enables VLMs to utilize visual trend cues and pattern recognition; and iv) a standardized instruction template specifying the desired style and scope of the output (see Appendix L.2). Crucially, the task is not limited to identifying trends, patterns, or summary statistics, but evaluates whether models can synthesize numeric evidence, visual cues, and semantic context into a coherent free-form natural language narrative that reflects not only what information is selected, but how it is composed and whether the resulting narrative is informative and correct, rather than producing templated or fragmentary observations.
Table 3: Evaluation results of generated captions against HR and synthetic (Synth) ground truths. The "Numeric Metrics" block comprises our tailored metrics (note: Numeric: numeric fidelity score, Mean Inf. = accuracy of detecting the time series mean). Bolded and underlined scores denote first and second places.
Model Lexical and Semantic Metrics Numeric Metrics
DeBERTa F1 SimCSE BLEU ROUGE-L METEOR Numeric Mean Inf. Max Inf. Min Inf.
HR Synth HR Synth HR Synth HR Synth HR Synth HR Synth HR Synth HR Synth HR Synth
Propr. Gemini 2.0 Flash 0.694 0.717 0.884 0.907 0.113 0.173 0.283 0.340 0.304 0.378 0.757 0.795 0.894 0.894 0.991 0.991 0.985 0.985
Claude 3 Haiku 0.676 0.697 0.856 0.873 0.086 0.137 0.261 0.305 0.286 0.349 0.669 0.712 0.842 0.841 0.983 0.983 0.981 0.980
GPT-4o 0.685 0.702 0.886 0.903 0.090 0.128 0.259 0.300 0.314 0.374 0.739 0.772 0.921 0.921 0.998 0.998 0.995 0.995
Pretrained InternVL 2.5 38b 0.682 0.704 0.867 0.889 0.085 0.144 0.262 0.314 0.294 0.367 0.740 0.790 0.865 0.865 0.983 0.982 0.978 0.978
LLaVA v1.6 0.641 0.655 0.796 0.814 0.064 0.088 0.234 0.267 0.248 0.290 0.597 0.618 0.488 0.489 0.792 0.791 0.614 0.614
Idefics 2 0.615 0.633 0.760 0.787 0.035 0.062 0.213 0.255 0.169 0.206 0.514 0.528 0.460 0.460 0.689 0.688 0.639 0.639
SmolVLM 0.572 0.586 0.691 0.709 0.029 0.053 0.195 0.230 0.169 0.203 0.508 0.540 0.476 0.494 0.860 0.860 0.647 0.647
QwenVL 0.638 0.656 0.809 0.832 0.060 0.095 0.229 0.268 0.239 0.282 0.544 0.564 0.496 0.496 0.788 0.789 0.601 0.601
QwenVL PAL 0.643 0.660 0.814 0.833 0.060 0.108 0.241 0.287 0.224 0.273 0.637 0.656 0.985 0.985 0.992 0.992 0.993 0.993
Llama 3.2 Vision 0.569 0.574 0.758 0.767 0.039 0.054 0.200 0.222 0.226 0.256 0.665 0.701 0.554 0.551 0.890 0.889 0.781 0.781
Gemma 3 27b 0.680 0.696 0.883 0.898 0.089 0.111 0.255 0.289 0.307 0.356 0.745 0.778 0.816 0.816 0.984 0.983 0.973 0.973
Phi-4 0.650 0.666 0.828 0.849 0.071 0.113 0.247 0.291 0.259 0.310 0.652 0.681 0.617 0.618 0.875 0.876 0.824 0.825
InternVL 2.5 8b 0.666 0.686 0.856 0.875 0.078 0.123 0.247 0.293 0.264 0.319 0.705 0.732 0.731 0.731 0.985 0.985 0.933 0.934
Gemma 3 12b 0.677 0.695 0.871 0.888 0.093 0.110 0.257 0.289 0.294 0.338 0.715 0.748 0.767 0.765 0.967 0.966 0.949 0.949
Finetuned LLaVA v1.6 0.644 0.659 0.802 0.819 0.064 0.090 0.234 0.269 0.250 0.292 0.589 0.613 0.500 0.498 0.827 0.826 0.615 0.615
Idefics 2 0.713 0.752 0.885 0.914 0.131 0.278 0.325 0.442 0.303 0.427 0.748 0.791 0.859 0.859 0.972 0.972 0.937 0.937
QwenVL 0.678 0.705 0.863 0.888 0.090 0.166 0.257 0.324 0.272 0.357 0.669 0.707 0.744 0.745 0.977 0.978 0.942 0.943
SmolVLM 0.601 0.616 0.751 0.771 0.056 0.105 0.224 0.273 0.197 0.243 0.604 0.632 0.433 0.437 0.796 0.796 0.606 0.606
Llama 3.2 Vision 0.672 0.686 0.851 0.867 0.087 0.130 0.266 0.309 0.295 0.352 0.740 0.761 0.711 0.710 0.942 0.942 0.877 0.876
Phi-4 0.664 0.682 0.862 0.880 0.078 0.117 0.243 0.285 0.290 0.349 0.703 0.743 0.651 0.651 0.853 0.855 0.817 0.817
3.5 Task 2: Diagnostic Q&A
To complement the generative evaluation, we introduce a suite of multiple-choice Q&A tasks designed to isolate and probe specific reasoning bottlenecks. These tasks are automatically derived from our base data triplets, with distractors designed to expose model failures in numeric precision and multimodal alignment. To ensure a high level of difficulty, we initially generated a pool of 44k questions per category and filtered them by removing those correctly answered by Qwen 2.5 Omni, focusing the benchmark on "hard" reasoning cases (see Appendix I.3 for an analysis of this filtering). The final diagnostic set consists of $910$ questions across the following tasks:
Time Series Matching. Models must identify the correct time series segment corresponding to a given caption from a set of distractors. These distractors are created via adversarial perturbations (shuffling, temporal reversal, Gaussian noise injection) to ensure that models cannot rely on simple matching and must instead understand the underlying temporal trend (Fig. 14; details in I.4).
Caption Matching. Given a numeric series, the model must select the correct caption from a set of semantic and numeric distractors. This isolates the model’s understanding of descriptive language from its ability to generate fluent text, specifically targeting sensitivities to perturbed statistical claims (Fig. 19; prompt in L.5).
Plot Matching. Model must select the correct line plot for a given caption and series. It specifically measures whether the visual modality provides a complementary signal or if the model remains biased toward textual priors (Fig. 20).
Comparative Reasoning. Given a pair of time series, the model must select the correct comparative statement (e.g., “Series A exhibits higher variance than Series B”) (Figs. 12131615). This probes the model’s ability to perform relative statistical assessment, a documented area of weakness for generative models (merrill2024language)
Temporal Matching. This task evaluates a model’s ability to perform exact temporal index alignment over time series data. In the first subtask (text-only), models are given a numeric time series as a sequence of values, along with a start timestamp and a query timestamp, and must select the correct value corresponding to the queried time. In the second subtask (vision-based), the numeric series is replaced by a line plot where only the starting time is labeled, requiring the model to infer temporal offsets visually before retrieving the correct value. These subtasks isolate discrete time indexing failures and test whether visual grounding helps or hinders temporal matching (Figs.17,18).
@@ 3.6 Evaluation Metrics
We evaluate models using a multi-faceted framework that goes beyond string matching to assess linguistic fluency, semantic alignment, and numeric grounding. Diagnostic Q&A tasks are evaluated using standard accuracy.
Linguistic and Semantic Alignment. We report BLEU (papineni2002bleu), ROUGE-L (chin2004rouge), and METEOR (banerjee2005meteor) to measure surface-level lexical overlap. To better capture temporal and semantic nuances, we additionally use embedding-based similarity metrics, namely DeBERTa Score (Zhang*2020BERTScore:) and SimCSE (gao2021simcse), which assess semantic consistency with human references independent of exact phrasing.
Numeric Grounding and Hallucination. The most critical challenge in TSC is ensuring that language is accurately grounded in numeric reality. Standard NLP metrics are notoriously "number-blind"; a model can achieve a high BLEU score while hallucinating incorrect values. To address this, we introduce two tailored metrics:
@@
    Numeric Fidelity Score: To measure the coverage of numeric information, we extract all numeric values from the HR reference (excluding dates) and match them with numbers in the model’s output. We calculate Accuracy (precision of the matched numbers) and Recall (the proportion of reference numbers successfully captured). The score is a weighted harmonic combination ($\lambda_{A}=0.3,\lambda_{R}=0.7$), reflecting our priority for recall; in a decision-making context, omitting a critical data point is often more detrimental than minor rounding differences.
@@
    Statistical Inference Accuracy: This set of metrics evaluates a model’s ability to correctly infer and verbalize key statistical references (mean, STD, min, max) from raw time series. We parse statistical claims from the generated text with an LLM and verify them against the true precomputed values (details in Appendix L.7). We report the percentage of correctly identified statistics within a $5\%$ relative error to factor out rounding errors (details on the tolerance design are in Appendix F.2). Crucially, this is a precision-oriented metric: models are not penalized for omitting a statistic but heavily penalized for reporting an incorrect one, thereby directly measuring numeric hallucination. (see Appendix F.3 for details on attempt analysis)
@@ 4 Experiments
We evaluate a number of VLMs on CaTS-Bench, covering both proprietary and open-source models. For TSC, we finetune some open-source models with the synthetic training data, and additionally consider a program-aided (PAL) model (gao2023PAL). Finetuning details can be found in Appendix D) All models are prompted with the same template-based format to ensure fair comparison, avoiding task- or architecture-specific prompt engineering. Model details and PAL are in Appendix E; the human baseline is in Appendix M.
4.1 Results on Time Series Captioning
@@
To ensure fair comparison across the domains, we report macro-averaged scores for each metric, mitigating sample size imbalances, as some domains contain more data, and preventing any domain from disproportionately influencing the results.
Our TSC results, shown in Table 3 , reveal several key insights. First, proprietary models consistently outperform pretrained open-source counterparts, with Gemini 2.0 Flash and GPT-4o establishing a strong baseline for high-quality caption generation. However, finetuned models demonstrate substantial improvements across nearly all metrics, with several open-source models narrowing the gap to proprietary baselines, showcasing the advantage of finetuning on our quality-verified synthetic data. Most notably, Idefics 2 stands out as the most effective learner: upon finetuning, it achieves the strongest performance among open-source models on several linguistic metrics and substantially narrows the gap to proprietary baselines, achieving, for instance, a DeBERTa $F_{1}$ score of $0.713$ and a ROUGE-L of $0.325$ against human ground truth. Although its gains in statistical inference are the most pronounced among finetuned models, it still lags behind the strongest proprietary systems in the Numeric category. In contrast, GPT-4o exhibits exceptional proficiency in statistical inference, followed closely by QwenVL PAL. The superior performance of the PAL variant over standard QwenVL highlights the substantial benefit of program-aided reasoning and code execution for accurate numerical tasks. Additionally, these results confirm that larger open-source models generally outperform their smaller counterparts.
Crucially, lexical and semantic scores are consistently lower when evaluated against HR ground truth than against synthetic ground truth. This consistent performance gap indicates that HR captions encode richer and less templatic reasoning patterns, and that the gains from finetuning primarily stem from exposure to higher-quality, more human-like data rather than increased data scale.
4.2 Results on Q&A Tasks
Figure 3 summarizes model performance on our Q&A tasks (see Table 19 for detailed results).